CAN SOMEONE FILL OUT THIS TABLE PLEASE?? I HAVE TRIED MULTIPLE TIME AND I AM NOT GETTING ANYTHING RIGHT.

2uF 10HZ 1kPREPARATION In an AC circuit capacitors and inductors have an effective resistance that restricts the flow of current, this iFor review, X_c (the capacitive resistance) is given by the equation 9-5 on page 30. Its 1 divided by 2 * P1 * the frequencyTable #1: Data Collection for RC circuit with AC input Frequency Calculated Calculated Z Measured (Hz) X. (1) (1) V. (V) Meas

THANK YOU :))))
Answer #1

ove will In filling up the table given forom ulas: use the following Xc 1. qT fc of is the frequeney C = 2 MF 2x 10-6.F. is tNow, VR = IR =( R > Z = R VR phone How, and VR are at difference The Vo = √2 + VR? so so Z= vuz + V VR R (measured)